Grade 11General Physics

Using the concept of relative motion answer the question. Velocity of the river with respect to ground is given by v. Width of river is d. A swimmer swims(wrt water) perpendicular to the current with acceleration a=2t where t is time starting from rest from the origin O at t=0.
Find the time of crossing the river.
Find the drift of the swimmer.

Rituraj Singh
12 Years agoGrade 11

2 Answers

Arun Kumar
12 Years ago
Hello Student,
a=2t
v=t^2 since at starting it must be zero
velocity of river=v
since the direction is perpendicular
so there is a drift distance.
dx/dx=t^2
d=t^3/3
=>t=(3d)^(1/3)
=>drift distance =v*(3d)^(1/3)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Abc
6 Years ago
a=2t
Since a= dv/dt
Substituting this in the equation 
dv/dt = 2t 
dv = 2t dt 
Integrating both sides 
∫dv = ∫2tdt 
v = 2
V is equal to dx/dt 
So , substituting this 
dx/dt = t 2
dx = t 2 dt
Integrating both sides we get 
x = t 3 /3 
here x is equal to width of river i.e d 
So , 
t = (3d) 1/3 
And since the velocity of river is v 
The drift will be vt 
= v (3d)1/3